solve
A defense attorney charges her client $4000 plus 120 per hour. The formula R=120n+4000 gives her revenue in dollars for n hours of work. What is her revenue for 100 hours of work? What is her revenue for 101 hours of work? By how much did the one extra hour of work increase the revenue?

Found 2 solutions by checkley75, funmath:
Answer by checkley75(3666)

R=120N+4000
R=120*100+4000
R=12000+4000
R=16000 FOR 100 HOURS OF WORK
R=120*101+4000
R=12120+4000
R=16120 FOR 101 HOURS WORKED OR
16120-16000=120 ADDITIONAL REVENUE FOR THE EXTRA HOUR.

Answer by funmath(2873)

solve
A defense attorney charges her client $4000 plus 120 per hour. The formula R=120n+4000 gives her revenue in dollars for n hours of work. (A) What is her revenue for 100 hours of work? (B) What is her revenue for 101 hours of work? (C)By how much did the one extra hour of work increase the revenue?
:(A)
Let n=100 and slove for R
R=120(100)+4000
R=12000+4000
R=$16000
: (B)
Let n=101 and solve for R
R=120(101)+4000
R=12120+4000
R=$16120
:(C)
R(101)-R(100)=16120-16000=$120
Happy Calculating!!!!